OpenStudy (sighn0more):

Can someone please check my answer to this logarithmic equation? Here is the question: log(base)2 (x + 1) + log(base)2 (x - 1) = log(base)2 8 My answer: x = 4.08 or -2.08

OpenStudy (sighn0more):

And if I got the answer wrong, what do I need to do to solve the problem correctly?

Nnesha (nnesha):

work > answer most important thin HOW did you get the answer not what is the answer :=)) so please explain how did you get that :=))

Nnesha (nnesha):

thing*

OpenStudy (sighn0more):

Okay! So I used the rules of logs to get: logbase2 (x + 1)(x-1) = logbase2 8, and then I expanded the binomial to get x^2 - 2x - 1 = logbase2 8. I then subtracted 8 from both sides, getting x^2 - 2x - 9 = 0, and then I used the quadratic equation to get my answer.

Nnesha (nnesha):

(x+1)(x-1) = ?? hint: difference of squares (a+b)(a-b) =a^2 -b^2 both parentheses are same but with opposite sign you can apply difference of squares

OpenStudy (sighn0more):

OH, I see. I completely forgot about that property. Okay, so it would be x^2 - 1^2

Nnesha (nnesha):

right which is same as x^2 -1 :=))

OpenStudy (sighn0more):

Awesome! Would I then subtract 8 from both sides of the equation?

Nnesha (nnesha):

yes right

OpenStudy (sighn0more):

Okay, cool! My new answer is x = 19.

Nnesha (nnesha):

hmm how ?

OpenStudy (sighn0more):

I put x^2 - 9 into the quadrtic equation > \[\sqrt{38}\] / 2 and got 19

OpenStudy (anonymous):

i wouldn't subtract 8 from both sides, i would add 1 to both sides and solve \[x^2=9\] in my head

Nnesha (nnesha):

well difference of squares x^2 -9= ??

OpenStudy (anonymous):

think of a number whose square is 9

OpenStudy (anonymous):

bet you get it on the first try

OpenStudy (sighn0more):

Ah! So it would be 3.

OpenStudy (sighn0more):

Or would it be both +3 and -3...

Nnesha (nnesha):

right \[x^2-9 =(x+3)(x-3)\] solve both binomial for x x+3 = 0 and x-3=0 x=-3 x=3 -3 is an extraneous solution so that wouldn't be the asnwer

Nnesha (nnesha):

when u substitute x for - 3 you will not get the equal sides that means -3 isn't a solution \[\log_2 (-3+1) +\log_2 (-3-1) \cancel{=} \log_2 8\]

Nnesha (nnesha):

so -3 isn't a solution

Nnesha (nnesha):

make sense ?? :=))

OpenStudy (sighn0more):

Yes! Perfect sense. Wow, that took a lot of steps; thanks for hanging in there with me!

Nnesha (nnesha):

np \o^_^o/ good work keep it up!